OpenStudy (anonymous):

Find the limit of (cos(3x))^(1/x^2) while x->0

OpenStudy (zarkon):

\[(\cos(3x))^{1/x^2}=exp\left(\frac{\ln(\cos(3x))}{x^2}\right)\]